# Beyond the Kolmogorov Barrier: A Learnable Weighted Hybrid Autoencoder for Model Order Reduction

### v3

@@ 3.3 Surrogate Modeling Datasets
3.3.1 1D Viscous Burgers’
One dimensional Viscous Burgers’ equation is represented by the following partial differential equation.
@@
$u(x,0)=u_{0},\quad x\in[0,L],\quad u(0,t)=u(L,t)=0.$
The initial condition used is given by $u(x,0)=\frac{x}{1+\sqrt{\frac{1}{t_{0}}}\exp\left(Re\frac{x^{2}}{4}\right)}$. This form of the initial condition allows for an analytical solution of the form $u(x,t)=\frac{\frac{x}{t+1}}{1+\sqrt{\frac{t+1}{t_{0}}}\exp\left(Re\frac{x^{2}}{4(t+4)}\right)}$, where $t_{0}=e^{Re/8}$ and $Re=\frac{1}{\nu}$. Data generation is carried out by computing the analytical solution on a grid of 128 points for 100 timesteps with a terminal time of 2s. Model training utilizes trajectories of 19 different $Re$ values ranging from 100 to 1900 in steps of 100, with 100 time steps in each. For testing, we hold out trajectories of 13 different $Re$ values ranging from 50 to 2450 in steps of 200.
A fully connected network using the same architecture as outlined in section 3.1.1 is used for the NN based encoder and decoder. The dimensionality reduction framework was trained with a learning rate of $31e-$ 3, batch size of 32 for 500 epochs using the Adam optimizer. To enhance convergence, we used a cyclic learning rate scheduler, which dynamically adjusts the learning rate during training. A latent space dimension of 2 is chosen for this problem.
Following the training and convergence of the dimensionality reduction framework, the encoder is utilized to obtain the reduced-order representations. The resulting embeddings, denoted as $\mathbf{z}\in\mathbb{R}^{r}$, are subsequently augmented with the Reynolds number ($Re$), thereby extending the latent space dimension from $r=2$ to $r^{\prime}=3$. This enriched representation, $\tilde{\mathbf{z}}=[\mathbf{z},Re]$, serves as the input to the LSTM model, where the objective is to approximate the discrete-time mapping $\tilde{\mathbf{z}}_{i+1}=\mathcal{F}_{\theta}(\tilde{\mathbf{z}}_{i})$, parameterized by $\theta$ (LSTM weights and biases of the NN). At this stage, the emphasis shifts from state compression to modeling the underlying temporal dynamics in the reduced-order manifold. The LSTM is then trained to learn the evolution of $\tilde{\mathbf{z}}$, facilitating the construction of a surrogate model capable of forecasting trajectory evolution given initial conditions and governing parameters.
The LSTM consists of 2 hidden layers with 40 neurons. The model is trained with a batch size of 32 for 400 epochs with a learning rate of $10^{-3}$ using Adam optimizer along with a cyclic learning rate scheduler. Data scaling is not performed at any stage of the training. The task here is to learn a parametric surrogate model capable of capturing the system state evolution via a reduced representation.
3.3.2 2D Shallow Water
@@
$\frac{\partial(\rho v)}{\partial t}+\frac{\partial(\rho uv)}{\partial x}+\frac{\partial}{\partial y}\left(\rho v^{2}+\frac{1}{2}\rho g\eta^{2}\right)$ $=0.$ (12)
@@
The initial conditions refer to a Gaussian pulse in the fluid column height, where the parameters $x_{0}$ and $y_{0}$ denote the initial location of the Gaussian pulse. We generate 90 trajectories for training, each representing a different initial location of pulse with a timestep of 0.001 s. The simulation is run for a final time of 0.5 s with the states being saved every 5 timesteps, leaving us 100 timesteps in each trajectory. Other 10 trajectories are held for testing purposes.
@@
where $\mathbf{u}=(u,v,w)$ is the velocity field, $b$ is the advection parameter and $\nu$ is the diffusion parameter. Simulation data for the 3D Viscous Burgers’ equation is obtained from APEBench dataset [27]. The advection and diffusion parameters are set to a value of -1.5 and 1.5 respectively emulating a decaying dynamics. We generate the simulation data for 50 different trajectories, each with a different initial condition on a grid of size $32^{3}$ containing 101 temporal snapshots in each trajectory. For this dataset, we use the first 50 timesteps for training and the rest for testing purposes.
We use a deep convolutional autoencoder for the NN part of the framework with the same architecture as outlined in section 3.1.2. The dimensionality reduction framework was trained with a learning rate of 1e-4, batch size of 20 for 3000 epochs using Adam optimizer. We choose a six-dimensional latent space to represent the system. The full orders states are scaled to zero mean and unit variance prior to training.
For this problem, we do not augment the latent space with any control parameters prior to training the LSTM model. The LSTM consists of 3 hidden layers with 128 neurons. The learning hyperparameters are similar to the 2D Shallow Water case. In this case we evaluate the model’s ability to extrapolate in time using the latent representation, which can be challenging given the dynamics is decaying.
@@ 4.1 Dimensionality Reduction Performance
To demonstrate the effectiveness of the proposed approach, we compare our model against three other methods. POD, vanilla deep autoencoder (AE), simple hybrid autoencoder, as autoencoders for the two chaotic PDE datasets in sections 3.1.1 and 3.1.2 with varying rank $r$ and resolution $N$. Sixteen ensembles of the model are trained by setting different random seeds. From fig. 4, both of the two hybrid approaches perform better than AE and POD. However, simple approach does not maintain its convergence with increasing rank in contrast to our proposed approach. In addition, our approach shows an order-of-magnitude improvement in generalization as compared to any of the other methods. The simple and our learnable weighted hybrid approach give nearly the same test error at low ranks, but their gap increases multiple times with increasing rank. Surprisingly, such substantial improvement merely requires a negligible additional trainable parameters (i.e. $r+Q\ll N$), as shown in table 5. For this dataset the convergence of our approach follows a SVD-like convergence, while the simple approach has a behavior similar to AE.
@@
For the more challenging 3D HIT dataset, fig. 5 shows that our proposed approach continues to outperform the other three methods in terms of generalization, especially when the resolution increases (e.g., $32^{3},64^{3}$ as opposed to $16^{3}$). It is important to note that the simple approach shows little improvement over POD at resolution of $32^{3}$ or $64^{3}$ while our approach excels. Again, this highlights the important role of learnable weights in our hybrid approach. Additionally for this dataset, we study the impact of activation function on the reconstruction performance by utilizing ReLU activation function for the NN part of the auto-encoder. It is worth noting that the performance of both the AE and simple hybrid approaches exhibit noticeable variations when the activation function is changed. In contrast, our approach shows minimal changes in performance, demonstrating the robustness of the proposed framework. A further comparison among four models is shown in fig. 6, which visualizes the improvement in performance of our learnable weighted approach, especially on the small-scale turbulence. It is interesting to note that both POD and simple hybrid approach predicts a overly negative $u_{y}$ near the right middle. In addition, as shown in table 6, introducing such learnable weights has almost no overhead on the total number of parameters.
@@ 4.2 Koopman Forecasting Performance
We compare the temporal extrapolation performance of the different methods in conjunction with the Koopman forecast model. It can be seen from fig. 7, that our approach achieves an order of magnitude lower testing error in comparison to the other methods for the traveling wave case, with simple hybrid performing similarly to POD as seen in some of the previous examples. For cylinder flow, both the simple hybrid and the proposed framework produce an order of magnitude lower testing error compared to POD and AE, and our framework further reduces the error by a factor of 2 compared to the simple hybrid. To better understand the prediction from these frameworks, fig. 8 shows the snapshot of last 1000 timesteps of the traveling wave for the training and testing regime and the prediction from the different methods. Prediction from POD and Simple Hybrid shows stationary dynamics. AE predicts a superposition of two different waves. The prediction of the proposed approach matches very well with the ground-truth data. The POD and simple hybrid approaches appear to be constrained by the expressivity of the model, preventing them from identifying a low-dimensional embedding of $r=2$ that effectively represents the system. This is a known limitation of POD for such problems. In this case, the simple hybrid was observed to be dominated by the POD component, resulting in an expressivity similar to that of POD. Although AE outperforms POD and Simple hybrid , its performance was found to be constrained by the optimization process, often getting stuck in some local minima. Meanwhile, the proposed framework effectively captures the mapping to and from the latent space and mitigates the optimization challenges, since it is initialized with POD values at the start of training. The performance of the frameworks on the training data indicates that our proposed framework effectively learns both the spatial and temporal characteristics of the dataset. This capability extends to the testing regime, demonstrating that our model generalizes well. For the cylinder flow dataset fig. 9, the proposed framework demonstrates superior generalization capabilities compared to alternative methods. Although the AE and simple hybrid approaches achieve performance comparable to our framework in the training regime, they exhibit notable degradation when the dynamics are forecast beyond the training horizon. In contrast, our model maintains high-fidelity reconstructions. This further emphasizes the significant improvement that can be made through the introduction of a few useful weighting parameters.
@@ 4.3 Surrogate Modeling Performance
We evaluate POD, AE, Simple Hybrid, and the proposed Learnable Weighted Hybrid for time-dependent PDE surrogate modeling, where dimensionality reduction is combined with LSTM-based time series prediction to assess the impact of reduced latent representations on system dynamics forecasting. The mean test error $L^{2}$ for the different methods in the three PDEs considered for surrogate modeling is shown in fig. 10. The hatched portion denotes the error contribution from the time series modeling, and the solid region represents the error due to dimensionality reduction. In all the cases, it can be seen that our proposed approach maintains a superior performance in comparison to the other methods. We obtain almost an order-of-magnitude reduction in generalization error for the 1D VB. For 2D Shallow Water and 3D VB, the simple hybrid approach does not show much improvement over POD, while AE has the highest testing error.
@@
To further illustrate the effectiveness of the proposed approach, we compare the reconstructed solutions for the 1D VB problem at the final time of 2s for a $Re$ value of 2450 in fig. 11. In such high Reynolds number scenarios, the solution exhibits the formation and propagation of shocks over time, making accurate reduced-order modeling particularly challenging. It can be noted that methods like AE and Simple Hybrid exhibit high-frequency oscillations in space, particularly near the shock, whereas our approach maintains a smooth profile that closely conforms to the ground truth. POD solution although smooth but doesn’t capture the location and profile of the shock accurately. This demonstrates that our method is well-suited for handling convection dominated scenarios.
@@
Next, we analyze the solution at the final time of 0.5s for the 2D Shallow Water case in fig. 12. Our approach effectively captures both large-scale and small-scale flow features within the solution domain. In contrast, POD and the Simple Hybrid approach primarily recover only the large-scale structures, while AE reconstruction contains non-sharp features. Similar comparison for the 3D VB can be found in fig. 13. While the predictions from our framework do not perfectly match the ground truth, they exhibit superior physical consistency and align more closely with the true solution than other methods, despite using only six latent dimensions. The other methods tend to predict an overly positive value for the velocities and generates totally nonphysical flow features.
@@
Figure 13: Generalization performance comparison on 3D VB data for a latent space rank of 6. The figure depicts the solution state at the final time for a testing trajectory. Each column corresponds to a particular state variable being compared.
@@ 4.4 Noise Robustness
Recent studies [28] in the deep learning community show that the sharpness of the minima, which describes the sensitivity of model loss with respect to perturbations in the model parameters, is a promising quantity that correlates with the generalization performance of deep networks. Let $\mathcal{D}_{\textrm{train}}=\left\{(s_{1},g_{1}),\ldots,(s_{n},g_{n})\right\}$ be the training data, that is, the set of features $s$ and target $g$ pairs, and $\ell_{i}(\mathbf{\Theta})$ be the loss of an NN model parametrized by weights $\mathbf{\Theta}$ and evaluated at the $i$th training sample point $\left(s_{i},g_{i}\right)$. Afterwards, the sharpness on a set of points $\mathcal{D}\subseteq\mathcal{D}_{\text{train}}$ can be defined as [29]: $\mathcal{S}(\mathbf{\Theta};\mathcal{D})\triangleq\max_{\|\delta\|_{2}\leq\rho}\frac{1}{|\mathcal{D}|}\sum_{i:\left(s_{i},g_{i}\right)\in\mathcal{D}}(\ell_{i}(\mathbf{\Theta}+\delta)-\ell_{i}(\mathbf{\Theta}))$ where $\delta$ is the perturbation introduced on the trainable parameters $\mathbf{\Theta}$ and $\rho$ refers to the perturbation radius. For example, table 4 shows the sharpness and reconstruction error for a particular training instance of 3D HIT data with grid points $32^{3}$ and rank 5. Our proposed framework has 1000 times less sharpness compared to the AE and simple hybrid approach. Since the sharpness of the minima is related to the resilience of the model in the presence of noisy data, we evaluate the reconstruction performance of AE, a learnable weighted simple hybrid framework under noisy testing data. Models trained on noise-free data are tested on data that contain random normal noise with zero mean and standard deviation equal to $10\%$, $20\%$ and $30\%$ of the maximum velocity magnitude superimposed over the flow field.
Method Sharpness Reconstruction $L^{2}$ Error Under Varying Noise Level (of Maximum Velocity)
@@
AE $16.69\pm 0.9$ 9.139.13{9.13}9.13 11.50 25.95 41.20
Simple Hybrid $22.61\pm 0.82$ 18.7618.76{18.76}18.76 22.26 36.93 58.12
Learnable Weighted Hybrid (Ours) $\mathbf{0.015}\pm\mathbf{0.001}$ $\mathbf{7.17}$ $\mathbf{8.12}$ $\mathbf{9.33}$ $\mathbf{9.43}$
@@
The reconstruction error for all models is shown in the same table. As the noise level increases, there is a drastic increase in the testing error for both the AE and the simple hybrid method, while our approach stays the same, demonstrating the robustness of this framework against noise in the unseen test data. A 2D slice of $u_{y}$ with and without 30% noise level is further visualized in fig. 14, which shows the improved robustness of our approach over other methods.

### v4

@@ 3.3 Surrogate Modeling Datasets
3.3.1 1D Viscous Burgers’ Equation
The one-dimensional Viscous Burgers’ equation is represented by the following partial differential equation.
@@
$u(x,0)=u_{0},\quad x\in[0,L],\quad u(0,t)=u(L,t)=0.$ (10)
The initial condition used is given by $u(x,0)=x/\left(1+\sqrt{\frac{1}{t_{0}}}\exp\left(Re\frac{x^{2}}{4}\right)\right)$. This initial condition allows for an analytical solution of the form $u(x,t)=\frac{x}{t+1}/\left({1+\sqrt{\frac{t+1}{t_{0}}}\exp\left(Re\frac{x^{2}}{4(t+4)}\right)}\right)$, where $t_{0}=e^{Re/8}$ and (Reynolds Number) $Re=\frac{1}{\nu}$. Data generation is carried out by computing the analytical solution on a grid of 128 points for 100 timesteps with a terminal time of 2s. Model training utilizes trajectories of 19 different $Re$ values ranging from 100 to 1900 in steps of 100, with 100 time steps in each. For testing, we hold out trajectories of 13 different $Re$ values ranging from 50 to 2450 in steps of 200.
A fully connected network using the same architecture as outlined in Section 3.1.1 is used for the NN based encoder and decoder. The dimensionality reduction framework was trained with a learning rate of $1\times 10{-3}$, batch size of 32 for 500 epochs using the Adam optimizer. To enhance convergence, we used a cyclic learning rate scheduler, which dynamically adjusts the learning rate during training. A latent space dimension of 2 is chosen for this problem.
Following the training and convergence of the dimensionality reduction framework, the encoder is utilized to obtain the reduced-order representations. The resulting embeddings, denoted as $\mathbf{z}\in\mathbb{R}^{r}$, are subsequently augmented with the $Re$ values, thereby extending the latent space dimension from $r=2$ to $r^{\prime}=3$. This enriched representation, $\tilde{\mathbf{z}}=[\mathbf{z},Re]$, serves as the input to the LSTM model, where the objective is to approximate the discrete-time mapping $\tilde{\mathbf{z}}_{i+1}=\mathcal{F}_{\theta}(\tilde{\mathbf{z}}_{i},Re)$, parameterized by $\theta$ (LSTM weights and biases of the NN). At this stage, the emphasis shifts from state compression to modeling the underlying temporal dynamics in the reduced-order manifold. The LSTM is then trained to learn the evolution of $\tilde{\mathbf{z}}$, facilitating the construction of a surrogate model capable of forecasting trajectory evolution given initial conditions and governing parameters.
The LSTM consists of 2 hidden layers with 40 neurons[24]. The model is trained with a batch size of 32 for 400 epochs with a learning rate of $10^{-3}$ using Adam optimizer along with a cyclic learning rate scheduler. Data scaling is not performed at any stage of the training. The task here is to learn a parametric surrogate model capable of capturing the system state evolution via a reduced representation.
3.3.2 2D Shallow Water Equations
@@
$\frac{\partial(\rho v)}{\partial t}+\frac{\partial(\rho uv)}{\partial x}+\frac{\partial}{\partial y}\left(\rho v^{2}+\frac{1}{2}\rho g\eta^{2}\right)$ $=0,$ (13)
@@
The initial conditions refer to a Gaussian pulse in the fluid column height, where the parameters $x_{0}$ and $y_{0}$ denote the initial location of the Gaussian pulse. We generate 90 trajectories for training, each representing a different initial location of pulse with a timestep of 0.001 s. The simulation is run for a final time of 0.5 s with the states being saved every 5 timesteps, leaving us 100 timesteps in each trajectory. Ten other trajectories are held for testing purposes.
@@
where $\mathbf{u}=(u,v,w)$ is the velocity field, $b$ is the advection parameter and $\nu$ is the diffusion parameter. Simulation data for the 3D Viscous Burgers’ equation is obtained from APEBench dataset [34]. The advection and diffusion parameters are set to a value of -1.5 and 1.5 respectively, emulating a decaying dynamics. We generate the simulation data for 50 different trajectories, each with a different initial condition on a grid of size $32^{3}$ containing 101 temporal snapshots in each trajectory. For this dataset, we use the first 50 timesteps for training and the rest for testing purposes.
We use a deep convolutional autoencoder for the NN part of the framework with the same architecture as outlined in Section 3.1.2. The dimensionality reduction framework was trained with a learning rate of $1\times 10^{-4}$, batch size of 20 for 3000 epochs using Adam optimizer. We choose a six-dimensional latent space to represent the system. The full orders states are scaled to zero mean and unit variance prior to training.
For this problem, we do not augment the latent space with any control parameters prior to training the LSTM model. The LSTM consists of 3 hidden layers with 128 neurons. The learning hyperparameters are similar to the 2D Shallow Water case. In this case we evaluate the model’s ability to extrapolate in time using the latent representation, which can be challenging given the dynamics are decaying.
@@ 4.1 Dimensionality Reduction Performance
To demonstrate the effectiveness of the proposed approach, we compare our model against three other methods. POD, vanilla deep autoencoder (AE), simple hybrid autoencoder, as autoencoders for the two chaotic PDE datasets in Sections 3.1.1 and 3.1.2 with varying rank $r$ and resolution $N$. Sixteen ensembles of the model are trained by setting different random seeds. From Figure 2, both of the two hybrid approaches perform better than AE and POD. However, simple hybrid approach does not maintain its convergence [^2] We use the term convergence in an empirical, asymptotic sense—referring to the monotonic decrease of training and generalization errors as the latent dimension $r$ increases. For sufficiently large $r$, the error tends toward a small bound $\epsilon$, exhibiting a qualitative decay resembling $\mathcal{O}(r^{-q})$, for some $q\in\mathbb{R}^{+}$. This is commonly desirable in reduced-order modeling, though perfect monotonicity may not hold—especially for regular autoencoders lacking optimal spectral alignment. Our learnable weighted model exhibits this trend and is bounded from below by POD, which often converges faster due to its optimality. Deviations may arise due to limited observability in coarse-grained turbulent fields. with increasing rank in contrast to our proposed approach. In addition, our approach shows an order-of-magnitude improvement in generalization as compared to any of the other methods. The simple and our learnable weighted hybrid approach give nearly the same test error at low ranks, but their gap increases multiple times with increasing rank. Surprisingly, such substantial improvement merely requires a negligible additional trainable parameters (i.e. $r+Q\ll N$), as shown in Table 6. For this dataset the convergence of our approach follows a SVD-like convergence, while the simple approach has a behavior similar to AE.
@@
For the more challenging 3D HIT dataset, Figure 3 shows that our proposed approach continues to outperform the other three methods in terms of generalization, especially when the resolution increases (e.g., $32^{3},64^{3}$ as opposed to $16^{3}$). It is important to note that the simple approach shows little improvement over POD at resolutions of $32^{3}$ or $64^{3}$ while our approach excels. Again, this highlights the important role of learnable weights in our hybrid approach. Additionally, for this dataset we study the impact of activation function on the reconstruction performance by utilizing ReLU activation function for the NN part of the auto-encoder. It is worth noting that the performance of both the AE and simple hybrid approaches exhibit noticeable variations when the activation function is changed. In contrast, our approach shows minimal changes in performance, demonstrating the robustness of the proposed framework. The latent space obtained using the proposed approach is also stable and reproducible (Appendix E).
To highlight the state-of-the-art performance of the proposed framework, we also perform a one-off comparison of its generalization ability on the complex 3D HIT case against the $\beta$-VAE baseline [18], as detailed in Appendix C. Our framework consistently achieves superior reconstruction performance across all grid resolutions and latent space ranks. Note that the improvement in performance is obtained at little to no additional computational overhead as shown in Table 7 in terms of the total number of training parameters and the training time per epoch which be seen in Figure 13 provided in Appendix B.
To better understand the relative roles of the linear POD basis and the nonlinear neural network in the proposed framework, we analyze their respective contributions to both the latent representation and the final reconstruction. A detailed breakdown of these contributions, across different resolutions and latent dimensions, is presented in Appendix D.
@@ 4.2 Koopman Forecasting Performance
We compare the temporal extrapolation performance of the different methods in conjunction with the Koopman forecast model. It can be seen from Figure 5, that our approach achieves an order of magnitude lower testing error in comparison to the other methods for the traveling wave case, with simple hybrid performing similarly to POD as seen in some of the previous examples. For cylinder flow, both the simple hybrid and the proposed framework produce an order of magnitude lower testing error compared to POD and AE, and our framework further reduces the error by a factor of 2 compared to the simple hybrid. To better understand the prediction from these frameworks, Figure 6 shows the snapshot of last 1000 timesteps of the traveling wave for the training and testing regime and the prediction from the different methods. Prediction from POD and Simple Hybrid show stationary dynamics. AE predicts a superposition of two different waves. The prediction of the proposed approach matches very well with the ground-truth data. The POD and simple hybrid approaches appear to be constrained by the expressivity of the model, preventing them from identifying a low-dimensional embedding of $r=2$ that effectively represents the system. This is a known limitation of POD for such problems. In this case, the simple hybrid was observed to be dominated by the POD component, resulting in an expressivity similar to that of POD. Although AE outperforms POD and Simple hybrid , its performance was found to be constrained by the optimization process, often getting stuck in some local minima. Meanwhile, the proposed framework effectively captures the mapping to and from the latent space and mitigates the optimization challenges, since it is initialized with POD values at the start of training. The performance of the frameworks on the training data indicates that our proposed framework effectively learns both the spatial and temporal characteristics of the dataset. This capability extends to the testing regime, demonstrating that our model generalizes well. For the cylinder flow dataset Figure 7, the proposed framework demonstrates superior generalization capabilities compared to alternative methods. Although the AE and simple hybrid approaches achieve performance comparable to our framework in the training regime, they exhibit notable degradation when the dynamics are forecast beyond the training horizon. In contrast, our model maintains high-fidelity reconstructions. This further emphasizes the significant improvement that can be made through the introduction of a few useful weighting parameters.
@@ 4.3 Surrogate Modeling Performance
We evaluate POD, AE, Simple Hybrid, and the proposed Learnable Weighted Hybrid for time-dependent PDE surrogate modeling, where dimensionality reduction is combined with LSTM-based time series prediction to assess the impact of reduced latent representations on system dynamics forecasting. The mean test error $L^{2}$ for the different methods in the three PDEs considered for surrogate modeling is shown in Figure 8. The hatched portion denotes the error contribution from the time series modeling, and the solid region represents the error due to dimensionality reduction. In all the cases, it can be seen that our proposed approach maintains a superior performance in comparison to the other methods. We obtain almost an order-of-magnitude reduction in generalization error for the 1D Viscous Burgers. For 2D Shallow Water and 3D Viscous Burgers, the simple hybrid approach does not show much improvement over POD, while AE has the highest testing error.
@@
To further illustrate the effectiveness of the proposed approach, we compare the reconstructed solutions for the 1D Viscous Burgers problem at the final time of 2s for a $Re$ value of 2450 in Figure 9. In such high Reynolds number scenarios, the solution exhibits the formation and propagation of shocks over time, making accurate reduced-order modeling particularly challenging. It can be noted that methods like AE and Simple Hybrid exhibit high-frequency oscillations in space, particularly near the shock, whereas our approach maintains a smooth profile that closely conforms to the ground truth. POD solution although smooth, does not capture the location and profile of the shock accurately. This demonstrates that our method is well-suited for handling convection dominated scenarios.
@@
Next, we analyze the solution at the final time of 0.5s for the 2D Shallow Water case in Figure 10. Our approach effectively captures both large-scale and small-scale flow features within the solution domain. In contrast, POD and the Simple Hybrid approach primarily recover only the large-scale structures, while AE reconstruction contains non-sharp features. Similar comparison for the 3D Viscous Burgers can be found in Figure 11. While the predictions from our framework do not perfectly match the ground truth, they exhibit superior physical consistency and align more closely with the true solution than other methods, despite using only six latent dimensions. The other methods tend to predict an overly positive value for the velocities and generates completely nonphysical flow features.
@@
Figure 11: Generalization performance comparison on 3D Viscous Burgers data for a latent space rank of 6. The figure depicts the solution state at the final time for a testing trajectory. Each column corresponds to a particular state variable being compared.
@@ 4.4 Noise Robustness
Recent studies [35] in the deep learning community show that the sharpness of the minima, which describes the sensitivity of model loss with respect to perturbations in the model parameters, is a promising quantity that correlates with the generalization performance of deep networks. Let $\mathcal{D}_{\textrm{train}}=\left\{(s_{1},g_{1}),\ldots,(s_{n},g_{n})\right\}$ be the training data, that is, the set of features $s$ and target $g$ pairs, and $\ell_{i}(\mathbf{\Theta})$ be the loss of an NN model parametrized by weights $\mathbf{\Theta}$ and evaluated at the $i$th training sample point $\left(s_{i},g_{i}\right)$. Afterwards, the sharpness on a set of points $\mathcal{D}\subseteq\mathcal{D}_{\text{train}}$ can be defined as [36]: $\mathcal{S}(\mathbf{\Theta};\mathcal{D})\triangleq\max_{\|\delta\|_{2}\leq\rho}\frac{1}{|\mathcal{D}|}\sum_{i:\left(s_{i},g_{i}\right)\in\mathcal{D}}(\ell_{i}(\mathbf{\Theta}+\delta)-\ell_{i}(\mathbf{\Theta}))$ where $\delta$ is the perturbation introduced on the trainable parameters $\mathbf{\Theta}$ and $\rho$ refers to the perturbation radius. For example, Table 4 shows the sharpness and reconstruction error for a particular training instance of 1D KS data with grid points $1024$ and rank 60 and Table 5 shows the sharpness and reconstruction error for a particular training instance of 3D HIT data with grid points $32^{3}$ and rank 5. The perturbation radius used here is 0.1, which is the largest value that can be applied without causing instability in the reconstructed outputs. Our proposed framework has 1000 times less sharpness compared to the AE and simple hybrid approach in the 3D HIT case. Since the sharpness of the minima is related to the resilience of the model in the presence of noisy data, we evaluate the reconstruction performance of AE, a learnable weighted simple hybrid framework under noisy testing data. Models trained on noise-free data are tested on data that contain random normal noise with zero mean and standard deviation equal to $10\%$, $20\%$ and $30\%$ of the maximum velocity magnitude superimposed over the flow field.
Method Sharpness
Reconstruction $L^{2}$ Error Under Varying
Noise Level (% of Maximum Velocity)
@@
AE
$2.21\times 10^{-6}$
$\pm 1.42\times 10^{-8}$
$5.97\times 10^{-03}$ $9.27\times 10^{-03}$ $4.89\times 10^{-02}$ $1.43\times 10^{-01}$
Simple Hybrid
$1.72\times 10^{-6}$
$\pm 6.07\times 10^{-7}$
$2.63\times 10^{-04}$ $6.78\times 10^{-04}$ $2.31\times 10^{-03}$ $5.63\times 10^{-02}$
Learnable Weighted
Hybrid (Ours)
$\mathbf{8.62\times 10^{-8}}$
$\pm\mathbf{9.47\times 10^{-10}}$
$\mathbf{2.56\times 10^{-05}}$ $\mathbf{3.12\times 10^{-05}}$ $\mathbf{6.78\times 10^{-05}}$ $\mathbf{1.18\times 10^{-04}}$
Table 4: Sharpness and reconstruction $L^{2}$ error for three deep autoencoders with rank $r=60$ on the 1D KS dataset at a resolution of 10241024 under varying noise levels in the test input.
Method Sharpness
Reconstruction $L^{2}$ Error Under Varying
Noise Level (% of Maximum Velocity)
No Noise 10% 20% 30%
AE $16.69\pm 0.9$ 9.13{9.13} 11.50 25.95 41.20
Simple Hybrid $22.61\pm 0.82$ 18.76{18.76} 22.26 36.93 58.12
Learnable Weighted
Hybrid (Ours)
$\mathbf{0.015}\pm\mathbf{0.001}$ $\mathbf{7.17}$ $\mathbf{8.12}$ $\mathbf{9.33}$ $\mathbf{9.43}$
@@
The reconstruction error for all models is shown in the same table. As the noise level increases, there is a drastic increase in the testing error for both the AE and the simple hybrid method, while our approach stays the same, demonstrating the robustness of this framework against noise in the unseen test data. For the 3D HIT case, a 2D slice of $u_{y}$ with and without 30% noise level is further visualized in Figure 12, which shows the improved robustness of our approach over other methods.